# SPAND: Sleep Prediction Architecture using Network Dynamics

### 20 Jan 2024 v1

@@ 8.1. Unlocking the Potential of Contagion
While existing literature has provided evidence for the existence of contagion through statistical testing (Udry, 2003), there exists a notable gap in the field: a lack of quantitative modeling that hinders its seamless integration into learning systems. Our proposed models have the capability to discern and incorporate underlying contagious behaviors into the learning of health labels. This capability proves especially valuable given the inherent scarcity of health and wellbeing labels, such as user emotions, in contrast to the high-resolution features readily obtainable through wearables and phones. This sparsity often necessitates some kind of downsampling of features, resulting in information loss. Because of the same limitation in our dataset, we were constrained to extracting statistical descriptions from multiple features rather than utilizing raw data. Leveraging contagion through graph networks offers a novel approach to harnessing the potential of wearable and smartphone data. This approach provides complementary information to the model, effectively reducing the loss of valuable information.
@@ 8.2. Enhancing sleep duration prediction through data aggregation
While comparing our model’s performance to that of using only time-series data in Table 2, we observe an improvement when the model aggregates data from multiple participants in a certain pattern, e.g. when employing CNN. However, the extent of this performance gain is limited due to the absence of information about the connections between these participants. Thus, the inclusion of graph networks significantly enhances the model’s predictive capabilities, providing indirect evidence of contagion effects. Nonetheless, it is crucial to acknowledge that not all friends or neighbors contribute equally, nor do they exert an equal magnitude of contagion. Therefore, when the model focuses its attention solely on important nodes using GAN, we not only witness performance improvements but also a substantial increase in model robustness as shown in Fig. 6, 8 and 11.
@@ 8.3. Robustness and node influence
Further robustness analysis highlighted that graph integration helps make the model more robust to data perturbations only when attention is paid to important nodes. Relying solely on the graph network for combining information from direct or indirect neighbors can make some nodes more vulnerable to noisy or missing data. Interestingly, the impact of perturbations on the performance varies depending on a node’s connectivity. Nodes directly connected to a substantial number of peers experience a less pronounced decline in performance compared to those linked to a select few highly influential nodes. This phenomenon is visually represented in Fig. 7 and 8. Firstly, nodes with higher degrees exhibit lower accuracy relative to their counterparts with fewer connections. The attention mechanism helps mitigate this effect to some extent and also makes the model more robust to perturbations. However, the users with high eigenvalue centrality experience a large drop in performance even with GAN. Thus, not all neighbors are equal, and different friends impact a node’s performance differently based on their own influence in the network.
@@ 8.4. Optimizing network size
Network size plays a pivotal role for two reasons: firstly, we want to capture all nodes that might have a contagious behavior leading to a certain user’s behavior. Therefore, we vary the graph size from 5 to 20 and observe the performance in Fig. 5(a). Our findings revealed a performance plateau at N=15. Larger graphs tended to introduce irrelevant data aggregation, although this could be partially alleviated by the implementation of an attention mechanism. This tradeoff can be optimized by carefully designing network size which is customized based on the user’s sociality, demographics, and lifestyle. For instance, in our study, all participants were university students who knew other people in the study; for these participants, we expect interactions with a larger number of fellow students. However, for specialized groups like caregivers, who may have extensive interactions within smaller circles (mainly care receivers), a smaller network size may be required for reliable performance. Graph-based models tend to be more sensitive to perturbations in neighboring nodes’ features due to their aggregation mechanism. Our observations in Fig. 6 and 12 indicated that in a graph of size 15, GCN remained robust even when up to one-third of nodes or when up to 1 day of the sequence were perturbed. However, as more users experienced perturbations, performance exhibited a linear decline. Notably, the CNN model displayed the most unstable performance and suffered the most significant drop when subjected to perturbations.
@@ 8.5. Harnessing the power of attention
While constructing social networks is relatively straightforward, their effective use in health label prediction presents multiple challenges. This is exemplified in Figures 7 and 13, where GCN performs poorly for nodes with a large number of direct neighbors that are also more susceptible to perturbations. However, the introduction of attention mechanisms significantly mitigates irrelevant data aggregation, resulting in improved performance. This underscores the critical role of attention in our approach.
The attention mechanism used in this work is based on feature correlation. However, it can be further improved by utilizing other information such as pre and post-surveys. The attention mechanism can also be customized based on network topologies. For instance, currently, attention weight is normalized by the weight of all neighbors. We noticed in our analysis that nodes with large eigenvalue centrality are more prone to noise. This insight can be used to do a weighted normalization or thresholding of the maximum number of important nodes when the node degree is very large.
8.6. Feature Importance
Figure 9. Feature Importance for different modalities for N=10 and L=3 days
SC, SCL, and SD stands for Skin Conductance, Skin Conductance Level, and Standard Deviation respectively. For more details on features, please refer to Tables. 4,5,6.
8.7. Impact of temporal module
@@ 8.8. Limitations
The dataset collected for this study was a human-centric dataset posing challenges in terms of the limited number of participants. This limitation makes it impossible to capture the entire social network of every participant (e.g., friends, family, coworkers) in one study. Therefore, the maximum graph size in our evaluation is 20 participants, which is not very large and does not represent very complex graph typologies. However, despite the limitations of the dataset, the architecture presented in this work shows promising results and highlights the potential that the integration of social networks can have in digital health systems. While this work focuses only on sleep duration, it can be easily extended to other health and well-being labels such as mood, emotion, and eating habits. Another limitation of this dataset is that all the participants were young adults at one university which results in sampling bias. In future works, the system needs to be tested with other diverse groups of the population. Moreover, the acquisition of ground truth involved user questionnaires, a method susceptible to subjective biases including recall limitations. These inherent biases can be mitigated in future works through the integration of sleep-tracking technologies(Rahman et al., 2015), such as smartwatches and under-mattress sleep trackers. The study was limited to metadata about calls, SMS, and screen on/off only. In future works, data from messenger apps or social media apps can be utilized to further expand the social networks used in the model.
Another limitation of the proposed system is that it is completely centralized. Information from all users is directed to one centralized model for making predictions. In order to deploy this system in large-scale real-life networks, this issue should be resolved through an elaborate distributive computing solution(Witt et al., 2022). Furthermore, the robustness analysis makes assumptions about the distribution of features in order to perturb them. We also used similar assumptions for standardization and removing outliers in data preprocessing. In future works, this assumption can be relaxed and domain knowledge about feature distribution can be utilized to further improve this analysis.
@@ 9. Conclusion
In this work, social networks are integrated with ubiquitous mobile and wearable sensor data to predict next-day sleep duration labels ( i.e., more or less than 8 hours) for multi-user networks. To achieve this goal, an architecture based on graph convolution networks is developed that aggregates information from different parts of the network. The proposed architecture showed distinct gains in prediction performance, corroborating the existence of contagion in sleep behavior and highlighting the potential of social networks to complement pervasive sensing data in mobile health systems. Further, robustness analysis exhibited the stability of the proposed architecture to perturbations in data. It also highlighted the significance of network topology as nodes with lower eigenvalue centrality are less impacted by disturbances compared to those with higher connectivity to influential nodes. This work provides a practical and scalable approach, to identifying dynamic social networks and integrating them into prediction models for health labels.
@@ B.1. Centrality Metrics
  • Eigenvalue centrality quantifies the influence of a node in a network by measuring the node’s closeness to influential parts of a network. It combines the degree of a node with the degree of its neighbors. For a graph $\mathcal{G}$ with adjacency matrix $A$, the eigenvalue centrality $C_{e}$ of a node $v$ is calculated by(Newman, 2010),
@@ Appendix C Additional Results
C.1. Statistical Analysis
The p-values obtained from the Tukey HSD test for pairwise comparison between proposed and benchmark models are reported in Table. 3
Proposed Model Benchmark p-value (Random Split) p-value (LOCO)
GCN-LSTM LSTM 0.0001 0.0000
GCN-LSTM CONV-LSTM 0.0002 0.001
GAN-LSTM LSTM 0.0000 0.0000
GAN-LSTM CONV-LSTM 0.0000 0.0000
Table 3. Tukey-HSD adjusted p-values.
C.2. Impact of Network Characteristics
@@
C.3. Robustness Analysis
Figure 11. Impact on MAPE by the varying quantity of perturbations in graphs of size 15 and sequence length of 3 days.
C.3.1. Perturbations in temporal domain
@@
Appendix D Dataset
D.1. Dataset Description
@@
D.1.1. Feature Description
There are four types of input data sources: wearable, mobile phone, questionnaires, and weather information. A detailed description of processed features obtained from these four sources is provided below.
  • Wearable device: The device worn by the participants on their wrist collected Skin Conductance (SC), Skin Temperature (ST), and 3-axis Accelerometer at 8 GHz. The units of Electro-Dermal Activity (EDA), ST, and Accelerometer were microsiemens ($\mu S$), Celsius (${}^{\circ}C$), and gravity ($G$) respectively. EDA is a physiological measurement that assesses the electrical conductance of the skin in response to sympathetic nervous system activity(Boucsein, 2012). EDA consists of Skin Conductance Level(SCL) and Skin Conductance(SC) Peaks (skin conductance response). SCL is the tonic electrical conductivity of skin while the peaks are phasic, short-term fluctuations in EDA due to a stimulus. In response to a stimulus, the EDA rises rapidly and then decays exponentially. These peaks are extracted using the procedure defined in (Taylor et al., 2015). The data for one 24-hour day were divided into three 7-hour and one 3-hour windows from 3-10H, 10-17H, 17-24H, and 0-3H. These time windows were chosen based on participants’ sleep routines. The average bedtime was 2:44 am with a standard deviation of 112 minutes. The average waketime was 9:30 am with a standard deviation of 114 minutes. For each window, multiple features from SC peaks, Skin Conductance Level (SCL), accelerometer, and temperature were computed. For SC Peak features, first peaks and artifacts were detected. Then, features listed in Table 4 were computed for both, all detected peaks, and only non-artifact peaks.
@@ D.1.1. Feature Description
  • Weather: Multiple features about the weather are extracted from DarkSky’s weather API (wea, 2016). The weather data contained information about temperature, sunlight, barometric pressure, wind, and the difference between that day’s weather and a rolling average. Details are provided in Table 6.
@@
Ground truth collection: We used as the ground truth sleep label value responses to the following questions that were asked in the morning survey:
@@
Finally time in bed, $y$ was calculated as follows,
$y=t_{2}-t_{s}-t_{1}-t_{4}$
@@
Duration Statistics For the duration of the call compute Mean, Median, Standard deviation and Total(Sum)
@@
D.2. Network Characteristics
D.2.1. Networks from survey information
Participants completed pre- and post-study surveys, ranking their connections (e.g., friends, roommates) from closest to farthest. For each question, participants could name up to three people in the order of how close they were to them. The first (closest) person was assigned 3 points, $2^{nd}$ closest person 2 points, and the $3^{rd}$ person 1 point. Then, points for all questions were added to get an aggregate score for closeness between two people. This score table was used to establish edge strengths in graphs. We created separate graphs for study participants from both surveys.
Many participants mentioned people who were not part of the study; we created separate pre-study and post-study graphs for only people who were study participants. When creating graphs between only-study participants, the percentage (rounded to the nearest integer) of unique connections that were retained from the network of all connections in the pre-study survey from cohort 1 to 7 was $18\%,36\%,17\%,15\%,21\%,10\%$ and $14\%$ respectively. For the post-study survey, the proportion retained from cohorts 1 to 7 was $15\%,34\%,13\%,17\%,20\%,10\%$, and $16\%$ respectively. Degree distributions in graphs with all participants were skewed due to out-of-study participants. Filtering to only-study participants provided a more balanced view, with most cohorts having a mean of approximately 2 nodes as shown in Table. 8.
To assess connection strength, we plotted edge strength distributions in Fig. 14. The boxplots showed substantial interquartile ranges within cohorts, indicating varying connection strengths. Out-of-study participants skewed graphs with all participants, resulting in numerous outliers. However, within-cohort participant graphs displayed significant variability in connection strength, suggesting that study participants had closer relationships with some individuals in the network. This evidence indicates that participants in the study knew each other (as required for study participation) and communicated via phone.
@@ D.2.1. Networks from survey information
D.2.2. Networks from Phone Data
We construct graphs from metadata collected from phones for each cohort. In order to understand the overall trend of how many users a node was connected to, the degree density for each cohort is presented in Fig. 15. For further understanding of the network characteristics within a cohort, the distribution of degree and edge strength is provided for SMS and call graphs in Fig. 16(a) and Fig. 16(b) respectively.

### 27 Jan 2024 v2

@@ 8.1. Unlocking the Potential of Contagion
While existing literature has provided evidence for contagion through statistical testing (Udry, 2003), there exists a notable gap in the field: a lack of quantitative modeling that hinders its seamless integration into learning systems. Our proposed models have the capability to discern and incorporate underlying contagious behaviors in the prediction task. This is especially valuable because of the inherent scarcity of health and wellbeing labels, such as user emotions, in contrast to the high-resolution features readily obtainable through wearables and phones. This sparsity often necessitates some kind of downsampling of features, resulting in information loss. Because of the same limitation in our dataset, we were constrained to extracting statistical descriptions from multiple features rather than utilizing raw data. Leveraging contagion through graph networks offers a novel approach to harnessing the potential of wearable and smartphone data. This approach provides complementary information to the model, effectively reducing the loss of valuable information.
@@ 8.2. Enhancing Sleep Duration Prediction through Data Aggregation
While comparing different models’ performance in Table 2, we observe an improvement when the model aggregates data from multiple participants in a certain pattern, e.g. when employing CNN. However, the extent of this gain is limited due to the absence of information about the connections between these participants. Thus, the inclusion of graph networks significantly enhances the model’s predictive capabilities, providing indirect evidence of contagion. Nonetheless, it is crucial to acknowledge that not all friends or neighbors contribute equally, nor do they exert an equal magnitude of contagion. Therefore, when the model focuses its attention solely on important nodes using GAN, we not only witness performance improvements but also a substantial increase in model robustness as shown in Fig. 7 and 9.
@@ 8.3. Robustness and Node Influence
Robustness analysis highlighted that graph integration helps make the model more robust to data perturbations only when attention is paid to important nodes. Relying solely on the graph neural network for combining information from other nodes can make some nodes more vulnerable to noisy or missing data. Interestingly, the impact of perturbations on the performance varies depending on a node’s connectivity. As shown in Fig. 8 and 9, nodes directly connected to a substantial number of peers experience a less pronounced decline in performance compared to those linked to a select few highly influential nodes. Firstly, nodes with higher degrees exhibit lower accuracy relative to their counterparts with fewer connections. The attention mechanism helps mitigate this effect to some extent and also makes the model more robust to perturbations. However, the users with high eigenvalue centrality experience a large drop in performance even with GAN. Thus, not all neighbors are equal, and different friends impact a node’s performance differently based on their own influence in the network.
@@ 8.4. Optimizing Network Size
Network size plays a pivotal role for two reasons: firstly, we want to capture all nodes that might have a contagious effect on user’s behavior. Therefore, we vary the graph size from 5 to 20 and observe the performance in Fig. 4(a). Our findings revealed a performance plateau at N=15. Larger graphs tended to introduce irrelevant data aggregation, although this could be partially alleviated by the implementation of an attention mechanism. This tradeoff can be optimized by carefully designing network size which is customized based on the user’s sociality, demographics, and lifestyle. For instance, in our study, all participants were university students who knew other people in the study; for these participants, we expect interactions with a larger number of fellow students. However, for specialized groups like caregivers, who may have extensive interactions within smaller circles (mainly care receivers), a smaller network size may be required for reliable performance. Graph-based models tend to be more sensitive to perturbations in neighboring nodes’ features due to their aggregation mechanism.
@@ 8.5. Harnessing the Power of Attention
While constructing social networks is relatively straightforward, their effective use in health label prediction presents multiple challenges. This is exemplified in Figs. 8 and 12, where GCN performs poorly for nodes with a large number of direct neighbors that are also more susceptible to perturbations. However, the introduction of attention mechanisms significantly mitigates irrelevant data aggregation, resulting in improved performance. This underscores the critical role of attention in our approach. The attention mechanism in this work is based on feature correlation. It can be further improved by utilizing complementary information from study surveys or customized based on network topologies. We noticed in our analysis that nodes with large eigenvalue centrality are more prone to noise. This insight can be used to do a weighted normalization of attention weight or thresholding the maximum number of important nodes when the node degree is very large.
@@ 8.6. Limitations
The collected dataset was a human-centric dataset posing challenges such as a limited number of participants. This limitation makes it impossible to capture the entire social network of each participant (e.g., friends, family, coworkers) in one study. Therefore, the maximum graph size in our evaluation is only 20, which does not represent very complex graph typologies. However, despite these limitations, the proposed architecture shows promising results and highlights the potential of social network integration in digital health systems. While this work focuses only on sleep duration, it can be easily extended to other health and well-being labels such as mood, emotion, and eating habits. Another limitation of this dataset is that all the participants were young adults at one university which results in sampling bias. In future works, the system needs to be tested with other diverse groups of the population. Moreover, the acquisition of ground truth involved user questionnaires, a method susceptible to subjective biases including recall limitations. These inherent biases can be mitigated in future works through the integration of sleep-tracking technologies (Rahman et al., 2015), such as smartwatches and under-mattress sleep trackers. The study was limited to metadata about calls, SMS, and screen on/off only. In future works, data from messenger or social media apps can be utilized to further expand the social networks used in the model.
Another limitation of this system is that it is completely centralized. Information from all users is directed to one centralized model for making predictions. In order to deploy this system in large-scale real-life networks, this issue should be resolved through an elaborate distributive computing solution(Witt et al., 2022). Furthermore, the robustness analysis makes assumptions about the distribution of features in order to perturb them. We also used similar assumptions for standardization and removing outliers in data preprocessing. In future works, this assumption can be relaxed and domain knowledge about feature distribution can be utilized to further improve this analysis.
@@ 9. Conclusion
In this work, social networks are integrated with ubiquitous mobile and wearable sensor data to predict next-day sleep duration labels ( i.e., more or less than 8 hours) for multi-user networks. To achieve this goal, an architecture based on graph convolution networks is developed that aggregates information from different parts of the network. The proposed architecture showed distinct gains in prediction performance, corroborating the existence of contagion in sleep behavior and highlighting the potential of social networks to complement pervasive sensing data in mobile health systems. Further, robustness analysis exhibited the stability of the proposed architecture to perturbations in data. It also highlighted the significance of network topology as nodes with lower eigenvalue centrality are less impacted by disturbances compared to those with higher connectivity to influential nodes. This work provides a practical and scalable approach, to identifying dynamic social networks and integrating them into prediction models for health labels. In addition to performance evaluation, a systematic analysis to characterize the system performance when deployed in the wild is also conducted. The robustness analysis exhibited the stability of the proposed architecture to perturbations in data. It also highlighted the significance of network topology as nodes with lower eigenvalue centrality are less impacted by disturbances compared to those with higher connectivity with influential nodes.
Acknowledgments
This work was supported by NSF (# 1840167 and # 2047296), NIH (R01GM10518), Samsung Electronics, and NEC Corporation. We thank SNAPSHOT study participants and collaborators.
@@ B.1. Centrality Metrics
  • Eigenvalue centrality quantifies the influence of a node in a network by measuring the node’s closeness to influential parts of a network. It combines the degree of a node with the degree of its neighbors. For a graph $\mathcal{G}$ with adjacency matrix $A$, the eigenvalue centrality $C_{e}$ of a node $v$ is calculated by (Arney, 2013),
@@ Appendix C Additional Results
C.1. Impact of Network Characteristics
@@
C.2. Robustness Analysis
C.2.1. Perturbations in temporal domain
@@
Appendix D Impact of temporal module
Appendix E Dataset
E.1. Dataset Description
@@
E.1.1. Feature Description
There are four types of data sources: wearable, mobile phone, questionnaires, and weather. A detailed description of processed features obtained from these four sources is provided below.
  • Wearable device: The device worn by the participants on their wrist collected Skin Conductance (SC), Skin Temperature (ST), and 3-axis Accelerometer at 8 GHz. The units of Electro-Dermal Activity (EDA), ST, and Accelerometer were microsiemens ($\mu S$), Celsius (${}^{\circ}C$), and gravity ($G$) respectively. EDA is a physiological measurement that assesses the electrical conductance of the skin in response to sympathetic nervous system activity(Boucsein, 2012). EDA consists of Skin Conductance Level(SCL) and Skin Conductance(SC) Peaks (skin conductance response). SCL is the tonic electrical conductivity of skin while the peaks are phasic, short-term fluctuations in EDA due to a stimulus. In response to a stimulus, the EDA rises rapidly and then decays exponentially. These peaks are extracted using the procedure defined in (Taylor et al., 2015). The data for one 24-hour day were divided into three 7-hour and one 3-hour windows from 3-10H, 10-17H, 17-24H, and 0-3H. These time windows were chosen based on participants’ sleep routines. The average bedtime and waketime were 2:44 am (±$\pm 112$ mins) and 9:30 am (±$\pm 114$ mins) respectively. For each window, multiple features from SC peaks, Skin Conductance Level (SCL), accelerometer, and temperature were computed. For SC Peak features, first peaks and artifacts were detected. Then, features listed in Table 3 were computed for both, all detected peaks, and only non-artifact peaks.
@@ E.1.1. Feature Description
  • Weather: Multiple features about the weather are extracted from DarkSky’s weather API (wea, 2016). Features include temperature, sunlight, barometric pressure, wind, and the difference between that day’s weather and a rolling average. Details are provided in Table 5.
@@
Ground truth collection: The following questions asked in the morning survey are used as ground truth sleep,
@@
Finally time in bed, $y$ was calculated as follows: $y=t_{2}-t_{s}-t_{1}-t_{4}$
@@
Duration Statistics For the duration of the call compute Mean, Median, Standard deviation, and Total(Sum)
@@
E.2. Network Characteristics
E.2.1. Networks from survey information
Participants ranked their connections (e.g., friends, roommates) in pre- and post-study surveys, assigning up to three people per question with decreasing points from 3 to 1 based on closeness. These rankings were aggregated into scores to determine edge strengths in graphs, with separate graphs created for study participants based on both surveys.
Many participants mentioned out-of-study people; we created separate pre-study and post-study graphs for in-study participants. When creating graphs between only-study participants, the percentage (rounded to the nearest integer) of unique connections that were retained from the network of all connections in the pre-study survey from cohort 1 to 7 was $18\%,36\%,17\%,15\%,21\%,10\%$ and $14\%$ respectively. For the post-study survey, the proportion retained from cohorts 1 to 7 was $15\%,34\%,13\%,17\%,20\%,10\%$, and $16\%$ respectively. Degree distributions in graphs with all participants were skewed due to out-of-study participants. Filtering to only-study participants provided a more balanced view, with most cohorts having a mean of approximately 2 nodes as shown in Table 7.
To assess connection strength, we plotted edge strength distributions in Fig. 13. The boxplots revealed large interquartile ranges within cohorts, suggesting diverse connection strengths. Out-of-study participants caused outliers in overall graphs, while in-cohort graphs showed notable variability, indicating closer relationships among study participants, confirming their communication and acquaintance as required for the study.
@@ E.2.1. Networks from survey information
E.2.2. Networks from Phone Data
We construct graphs from metadata collected from phones for each cohort. To understand how many users a node was connected to, the degree density for each cohort is presented in Fig. 14. For further understanding of the network characteristics within a cohort, the distribution of degree and edge strength is provided for SMS and call graphs in Fig. 15(a) and Fig. 15(b) respectively.